two angles are supplementary. One angle measures 30 less than twice the other. What are the measures of the angles?
Answer by ankor@dixie-net.com(22740)

"two angles are supplementary."
This means the sum of the two angle = 180 degrees
:
One angle measures 30 less than twice the other.
Let one angle = x
the other angle = (2x-30)
:
What are the measures of the angles?
x + (2x-30) = 180
3x = 180 + 30
3x = 210
x = 210/3
x = 70 degrees
:
The other angle = 2(70) - 30 = 110 degrees
:
70 + 110 = 180 as you would expect